Monetarists economists, like Milton Friedman, who believe in the Quantity theory of money, predict that an increase in money supply will increase

only  nominal interest rate
both  nominal interest rate and price level
only  price level
only real output
Answer #1

Answer

only price level

Clarification of why cash supply prompts inflation

Monetarists trust that in the transient speed (V) is settled This is on the grounds that the rate at which cash flows is dictated by institutional variables, for example how frequently laborers are paid does not change without a doubt. Milton Friedman let it out might shift a little however not without a doubt so it tends to be treated as settled

Monetarists likewise trust output Y is settled. They state it might differ in the short run yet not over the long haul (on the grounds that LRAS is inelastic and controlled by supply-side elements.)

In this way an expansion in the Money Supply will prompt an increment in inflation

Precedent 1

On the off chance that the absolute cash supply is at first £1000 and the speed of dissemination is 5.

The dimension of output (Y) is 5000 units.

£1000×5 = P (5000)

Thusly P = 1

On the off chance that the cash supply currently duplicates the condition =

2000×5 =P×5000

In this manner P = 2